# Structured coalescents, coagulation equations and multi-type branching processes

### v2

@@ 3.1. The action of the generator
To describe the action of the generator $A^{K}$ of the process of empirical measures ${\boldsymbol{\mu}}^{K}$ on functions in ${\mathcal{H}}$, we write
@@
where the operators $A^{K}_{M}$ and $A_{C}^{m}$ account for the contributions of migrations and coalescences, respectively. For the migration part, recall that any block-configuration ${\boldsymbol{c}}$ present in colony $i$ migrates to colony $j$ at rate $w_{i,j}K$, which has the effect on ${\boldsymbol{\mu}}^{K}$ of adding to coordinate $j$ a mass $1/K$ at ${\boldsymbol{c}}/s_{K}$ and removing from coordinate $i$ a mass $1/K$ at ${\boldsymbol{c}}/s_{K}$ (see Fig. 3 (right)). Since we are scaling time by $1/K$, we obtain
@@
3.2. Useful bounds
Lemma 3.1.
For any $t\geq 0$, we have
and
Proof.
$H^{F_{1},{\boldsymbol{f}}}({\boldsymbol{p}})=\sum_{i\in[d]}\langle{\boldsymbol{p}},\|\cdot\|_{1}^{2}\rangle.$
Recalling the definition of $\mu_{i}^{K}$ yields
$\frac{\mathrm{d}}{{\mathrm{d}t}}\psi_{1}(t)\leq b^{2}\alpha_{\mathrm{max}},$
Moreover, at time 00 we have
and (3.4) follows.
Moreover, at time 00, we have $\psi_{2}(0)=b^{4}/\gamma_{K}^{2}$. We conclude that
and (3.5) follows. ∎
@@ 3.3. Convergence of the generators
In this section, we will prove a uniform convergence result for the generator $A^{K}$. Since our final goal is to prove the convergence of the measures $({\boldsymbol{\mu}}^{K})_{K\geq 1}$ toward the solution of the coagulation equation (2.3) (resp. (2.5)), the generator of the limiting object should be given formally by
$\bar{A}G({\boldsymbol{p}})=\frac{{\rm d}}{{\rm d}t}G\circ u(\cdot,{\boldsymbol{p}})|_{t=0}$
where $u(\cdot,{\boldsymbol{p}})$ denotes the solution of Eq. (2.3) (resp. Eq. (2.5)) with initial value ${\boldsymbol{p}}$. For functions $G=H^{F,{\boldsymbol{f}}}\in{\mathcal{H}}$, a straightforward calculation shows that $\bar{A}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})=\bar{A}_{M}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})+\bar{A}_{C}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})\text{ with}$
@@
The result follows combining the bounds for the error arising from the migration and from the coalescent parts. ∎
@@ 3.4. Comparison results
In this section we will provide a comparison result between the total number of blocks in the multi-type coalescent ${\boldsymbol{\Pi}}^{K}$ and two Kingman coalescents. This result will allow us to get a hand on the order of magnitude of the colony sizes.
@@ 3.4.1. The coupling
Indeed, for $\ell\in{\mathbb{N}}_{0}^{d}$ with $|{\boldsymbol{\ell}}|-1\geq d$, Cauchy-Schwartz inequality yields
@@
Assume that we have constructed them up to the $k$-th transition (transitions refer to the times at which at least one of the three processes makes a jump; we consider time 00 as the 00-th transition), which brings $(L_{i}^{K})_{i\in[d]}$, ${\hat{L}}^{K}$ and $\tilde{L}^{K}$ respectively to states $\ell_{*},{\boldsymbol{\ell}},\ell^{*}$ satisfying $\ell_{*}\leq|{\boldsymbol{\ell}}|\leq\ell^{*}$ and $|{\boldsymbol{\ell}}|\geq d+1$. The next transition is then constructed as follows. We define
@@
and $C(\ell_{*},{\boldsymbol{\ell}},\ell^{*})\coloneqq\max\{\hat{c}(\ell_{*}),c({\boldsymbol{\ell}}),\tilde{c}(\ell^{*})\}$. We refer to $\hat{c}(\ell_{*}),c({\boldsymbol{\ell}})$ and $\tilde{c}(\ell^{*})$ as coalescence rates of $\ell_{*}$, ${\boldsymbol{\ell}}$, $\ell^{*}$, respectively. Then after an exponential time with parameter $\lambda(\ell_{*},{\boldsymbol{\ell}},\ell^{*})\coloneqq\rho({\boldsymbol{\ell}})+\max\{\hat{c}(\ell_{*}),c({\boldsymbol{\ell}}),\tilde{c}(\ell^{*})\}$
@@
    with probability $\rho({\boldsymbol{\ell}})/\lambda(\ell_{*},{\boldsymbol{\ell}},\ell^{*})$ a migration event takes place (affecting only ${\boldsymbol{\ell}}$). More precisely,
@@
    There are $6$ possible cases to consider that correspond to the different orderings of the coalescence rates of $\ell_{*},{\boldsymbol{\ell}},\ell^{*}$. We explain in detail one case; the others are analogous. We consider the case where $\hat{c}(\ell_{*})\geq c({\boldsymbol{\ell}})\geq\tilde{c}(\ell^{*})$. In this case, the transition
@@
    takes place. In addition, with probability $c({\boldsymbol{\ell}})/\hat{c}(\ell_{*})$ the state ${\boldsymbol{\ell}}$ transitions. The transition affects only one of its coordinates; it is the $i$-th coordinate with probability $\alpha_{i}\ell_{i}(\ell_{i}-1)/2c({\boldsymbol{\ell}})$, in which case the transition
@@
    takes place. Finally, only if ${\boldsymbol{\ell}}$ changes, the transition $\tilde{c}(\ell^{*})/c({\boldsymbol{\ell}})$
@@
As soon as $|{\boldsymbol{\ell}}|$ goes below $d+1$, we continue to carry out the coupling between $\hat{L}^{K}$ and $L^{K}$; the process $\tilde{L}^{K}$ can then be further constructed independently.
@@ 3.4.2. Kingman bounds
The following result on Kingman’s coalescent will be helpful in many proofs. We provide here a short proof for completeness.
@@
Let $(L_{\rho}^{K}(t))_{t\geq 0}$ be the block-counting process of a Kingman coalescent with merger rate $\rho$, which started with $N_{K}$ blocks at time 00. Then, for any $p\geq 1$, we have
@@
Dividing both sides by $v(s)^{1+1/p}$ and integrating both sides of the resulting inequality between 00 and $t$ and rearranging terms yields the result. ∎
@@ 4. Tightness and characterization of accumulation points
In this section we will prove the tightness of the sequence $\{{{\boldsymbol{\mu}}}^{K}\}_{K\geq 1}$ and characterize its accumulation points, paving the way to the desired convergence result.
@@ 4.1. Tightness
The next proposition states the tightness of $\{{\boldsymbol{\mu}}^{K}\}_{K\geq 1}$ and is the main result of this section.
@@
Let $t_{0}=0$ in the critical sampling case and $t_{0}=\varepsilon>0$ (with $\varepsilon>0$ fix, but arbitrary) in the large sampling case. For any $T>0$, the sequence of measure-valued processes $\left\{({\boldsymbol{\mu}}^{K}(t))_{t\in[t_{0},T]}\right\}_{K\geq 1}$ is tight in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d})^{d},w))$, where $(w)$ stands for the weak topology,
4.1.1. The proof strategy
To prove Proposition 4.1, we will show that the sequence of product measures
@@ 4.1.1. The proof strategy
is tight. The tightness of the sequence $\{{\boldsymbol{\mu}}^{K}\}_{K\geq 1}$ will then follow from the continuous mapping theorem (see [8, Thm. 2.7]).
@@
Our proof of the tightness of $\{{\mu_{\otimes}^{K}}\}_{K\geq 1}$ in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),w))$ follows a standard line of argumentation (see e.g. [16, 42, 41]), which, in our setting, amounts to proving that the following three conditions are met.
(P1) For every function $f_{\otimes}\in\mathcal{F}$, the sequence $\{(\langle\mu_{\otimes}^{K}(t),f_{\otimes}\rangle)_{t\in[t_{0},T]}\}_{K\geq 1}$ is tight in $D([t_{0},T],\mathbb{R})$.
(P2) $\limsup_{K\rightarrow\infty}\mathbb{E}\left[\sup_{t\in[t_{0},T]}\langle\mu_{\otimes}^{K}(t),\|\cdot\|_{2}^{2}\rangle\right]<\infty$.
(P3) Any accumulation point of $\left\{\mu_{\otimes}^{K}\right\}_{K\geq 1}$ in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),w))$ belongs to $C([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),w))$.
The strategy is analogous to the one used in [28, Thm. 7.4]), but they use a version of (P3) considering accumulation points in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),v))$, where $(v)$ stands for the vague topology. The fact that these three conditions imply the tightness of $\{{\mu_{\otimes}^{K}}\}_{K\geq 1}$ follows from [42, Thm. 1.1.8] (see Remark 4.2 below). According to Roelly’s criterium (see [39]), condition (P1) already implies tightness in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),v))$, where $(v)$ stands for the vague topology. Conditions (P2) and (P3) allow one to go from vague convergence to weak convergence (see [42, Lemma. 1.1.9]).
@@
Property (P2) slightly differs from the property stated in [42, Thm. 1.1.8], which reads as
$\lim_{k\rightarrow\infty}\limsup_{K\rightarrow\infty}\sup_{t\in[0,T]}\langle\mu_{\otimes}^{K}(t),\varphi_{k}(\|\cdot\|_{2})\rangle=0\text{, in probability,}$, (4.1)
where $\varphi_{k}$ is a function satisfying that $\mathds{1}_{\{\|{\boldsymbol{x}}\|_{2}\geq k\}}\leq\varphi_{k}(\|{\boldsymbol{x}}\|_{2})\leq\mathds{1}_{\{\|{\boldsymbol{x}}\|_{2}\geq k-1\}}.$ To see that (P2) is a stronger condition, note that Markov’s inequality implies that, for all $t\geq t_{0}$,
@@
Another application of Markov’s inequality shows that, under (P2), condition (4.1) holds.
@@
The proofs of (P1), (P2) and (P3) will be given in Lemmas 4.7,4.8 and 4.9, respectively.
4.1.2. A bound on the quadratic variations
In the next lemma, an important role is played by the functions ${\pi}_{I}:{\mathbb{R}}^{d}\to{\mathbb{R}}$, $I\subseteq[d]$, defined via ${\pi}_{I}({\boldsymbol{x}})=\prod_{j\in I}x_{j}.$
@@ 4.1.2. A bound on the quadratic variations
4.1.3. On the proof of (P1)
To prove (P1), we will use a classic result of Aldous and Rebolledo (see [1], [21]), which consists of two ingredients; the first is the content of the next lemma.
@@ 4.1.3. On the proof of (P1)
In the critical (resp. large) sampling case, for every fix $t\geq 0$ (resp. $t>0$), the sequence $\{\langle\mu_{\otimes}^{K}(t),f_{\otimes}\rangle\}_{K\geq 1}$ is tight.
@@
In particular, the two quantities are bounded from above by a function of $\delta$ that goes to 00 as $\delta\rightarrow 0$.
@@
uniformly in $K$, $\sigma$ and $\tau$. For the second expectation, we use Eq. (4.4) to find,
@@
Lemma 4.7 ((P1)).
Let $t_{0}=0$ in the critical sampling case and $t_{0}=\varepsilon>0$ in the large sampling case. For every function $f_{\otimes}\in\mathcal{F}$ and $t_{0}<T$, the sequence $\{(\langle{\mu_{\otimes}^{K}}(t),f_{\otimes}\rangle)_{t\in[0,T]}\}_{K\geq 1}$ is tight in $D([t_{0},T],\mathbb{R})$.
Proof.
The result follows as a consequence of Lemma 4.5 and Lemma 4.6 and the Aldous and Rebolledo criterium ([1], [21]). ∎
4.1.4. On the proof of (P2)
Now we proceed with the proof of condition (P2).
Lemma 4.8 ((P2)).
Proof.
@@ 4.1.4. On the proof of (P2)
$\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\langle\mu_{\otimes}^{K}(t),\|\cdot\|_{2}^{2}\rangle\bigg{]}$ $=\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{2}^{2}\rangle\prod_{n\in[d]\setminus\{i\}}\langle\mu_{n}^{K}(t),1\rangle\bigg{]}\leq\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{2}^{2}\rangle\bigg{]}.$
@@
$\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\langle\mu_{\otimes}^{K}(t),\|\cdot\|_{2}^{2}\rangle\bigg{]}\leq\mathbb{E}\bigg{[}\sum_{i\in[d]}\langle\mu_{i}^{K}(T),\|\cdot\|_{1}^{2}\rangle\bigg{]}\leq b^{2}\Big{(}\frac{1}{\gamma_{K}}+T\Big{)},$
@@
$\limsup_{K\rightarrow\infty}\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\langle\mu_{\otimes}^{K}(t),\|{\boldsymbol{x}}\|_{2}^{2}\rangle\bigg{]}\leq\limsup_{K\rightarrow\infty}b^{2}\Big{(}\frac{1}{\gamma_{K}}+T\Big{)}<\infty.$
@@
Moreover, according to Lemmas 3.3 and 3.4
@@
4.1.5. On the proofs of (P3) and Proposition 4.1
In this section we prove condition (P3) and conclude the proof of Proposition 4.1
Lemma 4.9 ((P3)).
Proof.
@@
We now have all the ingredients to prove the tightness of $\{{\mu_{\otimes}^{K}}\}_{K\geq 1}$.
Proof of Proposition 4.1.
Thanks to Lemmas 4.7, 4.8 and 4.9 we conclude that conditions (P1), (P2) and (P3) are satisfied. Hence the result follows from [42, Thm. 1.1.8]. ∎

### v3

@@ 3.1. The action of the generator
Remark 3.1.
In the critical sampling regime, the space $\mathbb{R}_{+}^{d}$ must be replaced by $\mathbb{N}_{0}^{d}$. In particular, the test functions $f_{i}$ then need only be bounded and measurable (no continuity assumptions are required on a discrete space), whereas the regularity assumptions on $F$ used for generator calculations apply in both regimes.
Throughout what follows (in particular, also in Section 4), definitions and arguments are presented in the large-sampling setting. The corresponding statements in the critical sampling regime are obtained by making the above distinction; with this modification, all definitions and arguments carry over.
Recall that the configuration of a block refers to the $d$-dimensional vector whose $i$-th coordinate, $i\in[d]$, records the number of atoms of color $i$. To describe the action of the generator $A^{K}$ of the empirical-measure process ${\boldsymbol{\mu}}^{K}$ on functions in $\mathcal{H}_{s}$, we decompose
@@
where $A^{K}_{M}$ and $A^{K}_{C}$ account for the contributions of migrations and coalescences, respectively. For the migration part, remember that any block configuration $\boldsymbol{c}$} present in colony $i$ migrates to colony $j$ at rate $w_{i,j}K$. This adds to the $j$-th coordinate of ${\boldsymbol{\mu}}^{K}$ a mass $1/K$ at ${\boldsymbol{c}}/s_{K}$ and removes the same mass from the $i$-th coordinate (see Fig. 4, right). Since time is rescaled by $1/K$, we obtain
@@ 3.2. Convergence of the generators
In this section, we prove a uniform convergence result for the generator $A^{K}$. Since our ultimate goal is to establish the convergence of the measures $({\boldsymbol{\mu}}^{K})_{K\geq 1}$ toward the solution of the coagulation equation (2.3) (resp. (2.5)), the generator of the limiting object should be given by
$\bar{A}G({\boldsymbol{p}})=\frac{{\rm d}}{{\rm d}t}\,G\circ u(\cdot,{\boldsymbol{p}})\big|_{t=0},$
where $u(\cdot,{\boldsymbol{p}})$ denotes the solution of Eq. (2.3) (resp. (2.5)) with initial value $\boldsymbol{p}$}, $G$ is a smooth enough function, and $\circ$ denotes composition of functions. A straightforward calculation shows that the operator $\bar{A}$ defined above acts on functions $G=H^{F,{\boldsymbol{f}}}\in{\mathcal{H}}$ via $\bar{A}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})=\bar{A}_{M}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}})+\bar{A}_{C}H^{F,{\boldsymbol{f}}}({\boldsymbol{p}}),$ with
@@
The result follows by combining the bounds for the errors arising from the migration and coalescence parts. ∎
3.3. Moment bounds
Lemma 3.3.
For any $t\geq 0$, we have
and
Proof.
$H^{F_{1},{\boldsymbol{f}}}({\boldsymbol{p}})=\sum_{i\in[d]}\langle p_{i},\|\cdot\|_{1}^{2}\rangle.$
Recalling the definition of $\mu_{i}^{K}$ yields
$\frac{\mathrm{d}}{{\mathrm{d}t}}\psi_{1}(t)\leq b^{2}\alpha_{\mathrm{max}},$
Moreover, at time 0 we have
and (3.5) follows.
Moreover, at time 0, we have $\psi_{2}(0)=b^{4}/\gamma_{K}^{2}$. We conclude that
and (3.6) follows. ∎
@@ 3.4. Comparison results
In this section we will provide a comparison result between the total number of blocks in the structured coalescent ${\boldsymbol{\Pi}}^{K}$ and two Kingman coalescents. This result will allow us to get a hand on the order of magnitude of the colony sizes.
@@ 3.4.1. The coupling
Indeed, for $\ell\in{\mathbb{N}}_{0}^{d}$ with $|{\boldsymbol{\ell}}|-1\geq d$, Cauchy-Schwarz inequality yields
@@
Assume that we have constructed them up to the $k$-th transition (transitions refer to the times at which at least one of the three processes makes a jump; we consider time 0 as the 0-th transition), which brings $(L_{i}^{K})_{i\in[d]}$, ${\hat{L}}^{K}$ and $\tilde{L}^{K}$ respectively to states $\ell_{*},{\boldsymbol{\ell}},\ell^{*}$ satisfying $\ell_{*}\leq|{\boldsymbol{\ell}}|\leq\ell^{*}$ and $|{\boldsymbol{\ell}}|\geq d+1$. The next transition is then constructed as follows. We define
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3.4.2. Kingman bounds
The following result on Kingman’s coalescent will be helpful in many proofs. We provide a short proof for completeness.
@@
Let $(L_{\rho}^{K}(t))_{t\geq 0}$ be the block-counting process of a Kingman coalescent with merger rate $\rho$, which started with $N_{K}$ blocks at time 0. Then, for any $p\geq 1$, we have
@@
Dividing both sides by $v(s)^{1+1/p}$ and integrating both sides of the resulting inequality between 0 and $t$ and rearranging terms yields the result. ∎
@@ 4. Tightness and characterization of accumulation points
Having proved convergence of the generators in Section 3.2, we now turn to the proof of tightness of the sequence $\{{\boldsymbol{\mu}}^{K}\}_{K\geq 1}$ (see Proposition 4.1 in Section 4.1) and to the characterization of its accumulation points (see Proposition 4.10 in Section 4.2), thereby paving the way to the desired convergence result.
@@ 4.1. Tightness
The next proposition establishes the announced tightness of the sequence $\{{\boldsymbol{\mu}}^{K}\}_{K\geq 1}$.
@@
Let $t_{0}=0$ in the critical sampling case and $t_{0}=\varepsilon>0$ (with $\varepsilon>0$ fixed, but arbitrary) in the large sampling case. For any $T>t_{0}$, the sequence of measure-valued processes $\left\{({\boldsymbol{\mu}}^{K}(t))_{t\in[t_{0},T]}\right\}_{K\geq 1}$ is tight in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d})^{d},w))$, where $(w)$ stands for the weak topology,
The proof of this result follows a classical strategy and relies on several key ingredients. For clarity, we first establish the intermediate results associated with these ingredients and explain how they combine to yield the desired result. The remainder of the section is then devoted to proving these intermediate results.
4.1.1. The (skeleton of the) proof of Proposition 4.1
To prove tightness of the sequence $\{{\boldsymbol{\mu}}^{K}\}_{K\geq 1}$, we will show that the sequence of product measures
@@
is tight in the weak topology. The tightness of $\{{\boldsymbol{\mu}}^{K}\}_{K\geq 1}$ follows then as an application of the continuous mapping theorem (see [8, Thm. 2.7]).
@@
To prove the tightness of $\{{\mu_{\otimes}^{K}}\}_{K\geq 1}$ in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),w))$, we follow a standard approach (see, e.g. [17, 44, 43]). In our setting, this approach amounts to establishing the following three lemmas.
Lemma 4.2 (Tightness of integrals).
Let $t_{0}=0$ in the critical sampling case and $t_{0}=\varepsilon>0$ in the large sampling case. For every function $f_{\otimes}\in\mathcal{F}$ and $T>t_{0}$, the sequence $\{(\langle{\mu_{\otimes}^{K}}(t),f_{\otimes}\rangle)_{t\in[t_{0},T]}\}_{K\geq 1}$ is tight in $D([t_{0},T],\mathbb{R})$.
Lemma 4.3 (Uniform-in-time moment bound).
Lemma 4.4 (Continuity of accumulation points).
Assuming these three lemmas, Proposition 4.1 follows directly from [44, Thm. 1.1.8].
This strategy is analogous to the one used in [30, Thm. 7.4], although the authors there employ a version of Lemma 4.4 formulated for accumulation points in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),v))$, where $v$ denotes the vague topology. According to Roelly’s criterion (see [41]), Lemma 4.2 alone already ensures tightness in $D([t_{0},T],(\mathcal{M}_{f}(\mathbb{R}_{+}^{d^{2}}),v))$. Lemmas 4.3 and 4.4 then allow one to upgrade vague convergence to weak convergence (see [44, Lemma 1.1.9]).
@@
The bound in Lemma 4.3 slightly differs from the property stated in [44, Thm. 1.1.8], which reads as
$\lim_{k\rightarrow\infty}\limsup_{K\rightarrow\infty}\sup_{t\in[0,T]}\langle\mu_{\otimes}^{K}(t),\varphi_{k}(\|\cdot\|_{2})\rangle=0\text{, in probability,}$ (4.1)
where $\varphi_{k}$ is a function satisfying $\mathds{1}_{\{\|{\boldsymbol{x}}\|_{2}\geq k\}}\leq\varphi_{k}(\|{\boldsymbol{x}}\|_{2})\leq\mathds{1}_{\{\|{\boldsymbol{x}}\|_{2}\geq k-1\}}.$ Note that Markov’s inequality implies that, for all $t\geq t_{0}$,
@@
Another application of Markov’s inequality shows that, under the bound in Lemma 4.3, condition (4.1) holds.
@@
The remainder of Section 4.1 is devoted to proving Lemmas 4.2, 4.3 and $4.$
4.1.2. On the proof of Lemma 4.2
The proof of Lemma 4.2 is based on a classical result of Aldous and Rebolledo (see [1, 22]), which relies on two main ingredients: (i) tightness of integrals of functions with respect to $\mu_{\otimes}^{K}$} at fixed times (see Lemma 4.8); and (ii) bounds on the martingale and finite-variation parts of $H({\boldsymbol{\mu}}^{K})$ for an appropriate class of test functions $H$ (see Lemma 4.9).
We begin with an intermediate result that provides bounds on the increments of the quadratic variation of $H({\boldsymbol{\mu}}^{K})$ and on $A^{K}H$ along ${\boldsymbol{\mu}}^{K}$ for a suitable class of test functions $H$ (see Lemma 4.7). This result will be crucial for establishing the first ingredient and will also be used in Section 4.2. To this end, we introduce the functions ${\pi}_{I}:{\mathbb{R}}^{d}\to{\mathbb{R}}$, for $I\subseteq[d]$, defined by ${\pi}_{I}({\boldsymbol{x}})=\prod_{j\in I}x_{j}.$
@@
Let us now turn to the first ingredient, which is the content of the next lemma.
@@
In the critical (resp. large) sampling case, for every fixed $t\geq 0$ (resp. $t>0$), the sequence $\{\langle\mu_{\otimes}^{K}(t),f_{\otimes}\rangle\}_{K\geq 1}$ is tight.
@@
In particular, the two quantities are bounded from above by a function of $\delta$ that goes to 0 as $\delta\rightarrow 0$.
@@
uniformly in $K$, $\sigma$ and $\tau$. For the second expectation, we use Eq. (4.4) to find
@@
We conclude this section with the proof of Lemma 4.2.
Proof of Lemma 4.2.
The result follows as a consequence of Lemma 4.8 and Lemma 4.9 and the Aldous–Rebolledo tightness criterium ([1], [22]). ∎
4.1.3. On the proof of Lemma 4.3
We now establish the announced uniform-in-time moment bounds.
Proof of Lemma 4.3.
@@
$\mathbb{E}\bigg[\sup_{t\in[0,T]}\langle\mu_{\otimes}^{K}(t),\|\cdot\|_{2}^{2}\rangle\bigg]$ $=\mathbb{E}\bigg[\sup_{t\in[0,T]}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{2}^{2}\rangle\prod_{n\in[d]\setminus\{i\}}\langle\mu_{n}^{K}(t),1\rangle\bigg]$
$\leq b^{d-1}\mathbb{E}\bigg[\sup_{t\in[0,T]}\sum_{i\in[d]}\langle\mu_{i}^{K}(t),\|\cdot\|_{2}^{2}\rangle\bigg].$
@@
$\mathbb{E}\bigg[\sup_{t\in[0,T]}\langle\mu_{\otimes}^{K}(t),\|\cdot\|_{2}^{2}\rangle\bigg]\leq b^{d-1}\mathbb{E}\bigg[\sum_{i\in[d]}\langle\mu_{i}^{K}(T),\|\cdot\|_{1}^{2}\rangle\bigg]\leq b^{d+1}\Big(\frac{1}{\gamma_{K}}+T\Big),$
@@
$\limsup_{K\rightarrow\infty}\mathbb{E}\bigg[\sup_{t\in[0,T]}\langle\mu_{\otimes}^{K}(t),\|{\boldsymbol{x}}\|_{2}^{2}\rangle\bigg]\leq\limsup_{K\rightarrow\infty}b^{d+1}\Big(\frac{1}{\gamma_{K}}+T\Big)<\infty.$
@@
Moreover, according to Lemmas 3.4 and 3.5
@@
4.1.4. On the proofs of Lemma 4.4
We now turn to establishing the continuity of the accumulation points of the sequence $\{{\mu_{\otimes}^{K}}\}_{K\geq 1}$.
Proof of Lemma 4.4.